Income Statement for a Manufacturing Company

Two items are omitted from each of the following three lists of cost of goods sold data from a manufacturing company income statement. Determine the amounts of the missing items, identifying them by letter.

Finished goods inventory, June 1 $51,900 $14,500 (e)
Cost of goods manufactured 275,100 (c) 50,800
Cost of finished goods available for sale (a) $137,800 $79,800
Finished goods inventory, June 30 63,400 28,900 (f)
Cost of goods sold (b) (d) $73,400
a. $
b. $
c. $
d. $
e. $
f. $
Answer #1

Finished goods inventory, June 1+Cost of goods manufactured=Cost of finished goods available for sale

Cost of finished goods available for sale-Finished goods inventory, June 30=Cost of goods sold

Finished goods inventory, June 1 51900 14500 (79800-50800)=$29,000
Cost of goods manufactured 275100 (137800-14500)=$123,300 50800
Cost of finished goods available for sale (51900+275100)=$327,000 137800 79800
Finished goods inventory, June 30 63400 28900 (79800-73400)=$6400
Cost of goods sold (327000-63400)=$263600 (137800-28900)=$108,900 73400.